# Pretraining Decision Transformers with Reward Prediction for In-Context Multi-task Structured Bandit Learning

### v1

@@ Abstract
In this paper, we study multi-task structured bandit problem where the goal is to learn a near-optimal algorithm that minimizes cumulative regret. The tasks share a common structure and the algorithm exploits the shared structure to minimize the cumulative regret for an unseen but related test task. We use a transformer as a decision-making algorithm to learn this shared structure so as to generalize to the test task. The prior work of pretrained decision transformers like DPT requires access to the optimal action during training which may be hard in several scenarios. Diverging from these works, our learning algorithm does not need the knowledge of optimal action per task during training but predicts a reward vector for each of the actions using only the observed offline data from the diverse training tasks. Finally, during inference time, it selects action using the reward predictions employing various exploration strategies in-context for an unseen test task. We show that our model outperforms other SOTA methods like DPT, and Algorithmic Distillation (AD) over a series of experiments on several structured bandit problems (linear, bilinear, latent, non-linear). Interestingly, we show that our algorithm, without the knowledge of the underlying problem structure, can learn a near-optimal policy in-context by leveraging the shared structure across diverse tasks. We further extend the field of pre-trained decision transformers by showing that they can leverage unseen tasks with new actions and still learn the underlying latent structure to derive a near-optimal policy. We validate this over several experiments to show that our proposed solution is very general and has wide applications to potentially emergent online and offline strategies at test time. Finally, we theoretically analyze the performance of our algorithm and obtain generalization bounds in the in-context multi-task learning setting.
@@ 1 Introduction
In this paper, we study multi-task bandit learning with the goal of learning an algorithm that discovers and exploits structure in a family of related tasks. In multi-task bandit learning, we have multiple distinct bandit tasks for which we want to learn a policy. Though distinct, the tasks share some structure, which we hope to leverage to speed up learning on new instances in this task family. Traditionally, the study of such structured bandit problems has relied on knowledge of the problem structure like linear bandits (Li et al., 2010; Abbasi-Yadkori et al., 2011; Degenne et al., 2020), bilinear bandits (Jun et al., 2019), hierarchical bandits (Hong et al., 2022a, b), Lipschitz bandits (Bubeck et al., 2008, 2011; Magureanu et al., 2014), other structured bandits settings (Riquelme et al., 2018; Lattimore and Szepesvári, 2019; Dong et al., 2021) and even linear and bilinear multi-task bandit settings (Yang et al., 2022a; Du et al., 2023; Mukherjee et al., 2023). When structure is unknown an alternative is to adopt sophisticated model classes, such as kernel machines or neural networks, exemplified by kernel or neural bandits (Valko et al., 2013; Chowdhury and Gopalan, 2017; Zhou et al., 2020; Dai et al., 2022). However, these approaches are also costly as they learn complex, nonlinear models from the ground up without any prior data (Justus et al., 2018; Zambaldi et al., 2018).
@@
In this paper, we consider an alternative approach of synthesizing a bandit algorithm from historical data where the data comes from recorded bandit interactions with past instances of our target task family. Concretely, we are given a set of state-action-reward tuples obtained by running some bandit algorithm in various instances from the task family. We then aim to learn a transformer (Vaswani et al., 2017) from this data that can in-context learn on new task instances. Laskin et al. (2022) consider a similar goal and introduce the Algorithm Distillation (AD) method, however, AD aims to copy the algorithm used in the historical data and thus is limited by the ability of the data collection algorithm. Lee et al. (2023) develop an approach, DPT, that enables learning a transformer that obtains lower regret in-context bandit learning compared to the algorithm used to produce the historical data. However, this approach requires knowledge of the optimal action at each stage of the decision process. In real problems, this assumption is hard to satisfy and we will show that DPT performs poorly when the optimal action is only approximately known. With this past work in mind, the goal of this paper is to answer the question:
@@
Can we learn an in-context bandit learning algorithm that obtains lower regret than the algorithm used to produce training data without knowledge of optimal actions in each training task?
@@
To answer this question, we introduce a new transformer methodology that obviates the need for knowledge of the optimal action in the in-context data –- a piece of information that is often inaccessible. Our key observation is that while the mean rewards of each action change from task to task, certain probabilistic dependencies are persistent across all tasks with a given structure (Yang et al., 2020, 2022a; Mukherjee et al., 2023). These probabilistic dependencies can be learned from the pretraining data and exploited to better estimate mean rewards and improve performance in a new unknown test task. The nature of the probabilistic dependencies depends on the specific structure of the bandit and can be complex (i.e., higher-order dependencies beyond simple correlations). We propose to use transformer models as a general-purpose architecture to capture the unknown dependencies by training transformers to predict the mean rewards in each of the given trajectories (Mirchandani et al., 2023; Zhao et al., 2023). The key idea is that transformers have the capacity to discover and exploit complex dependencies in order to predict the rewards of all possible actions in each task from a small history of action-reward pairs in a new task. This paper demonstrates how such an approach can achieve state-of-the-art performance, relying solely on historical data, without the need for any supplementary information like the action features or knowledge of the complex reward models. Our new decision transformer learns to adapt, in-context, to novel actions as well as new tasks.
@@
    We introduce a new training procedure where predicting the next reward for all arms circumvents the issue of requiring access to the optimal (or approximately optimal) action during training time.
@@
    We demonstrate empirically that this training procedure results in lower regret in a wide series of tasks (such as linear, nonlinear, bilinear, and latent bandits) compared to DPT, AD (and other in-context decision-making algorithms).
@@
    We also show that our training procedure leverages the latent structure even when new actions are introduced both during training and testing time.
@@ 2 Background
In this section, we first introduce our notation and the multi-task, structured bandit setting. We then formalize the in-context learning model studied in Laskin et al. (2022); Lee et al. (2023); Sinii et al. (2023); Lin et al. (2023); Ma et al. (2023); Liu et al. (2023c, a). We choose DPT (Lee et al., 2023) as a representative example of this learning model and explain the DPT model for the online setting.
@@ 2.1 Preliminaries
In this paper, we consider the multi-task linear bandit setting (Du et al., 2023; Yang et al., 2020, 2022a). In the multi-task setting, we denote each task as $m$. Each task $m$ constitutes of the action set $|\mathcal{A}|=A$ and the feature space $\mathcal{X}$. The actions in $\mathcal{A}$ are indexed by $a=1,2,\ldots,A$. The feature of each action is denoted by $\mathbf{x}(m,a)\in\mathbb{R}^{d}$ and $d\ll A$. Denote $\triangle(\mathcal{A})$ as the probability simplex over the action space $\mathcal{A}$. We denote a policy as $\pi:\mathcal{A}\rightarrow[0,1]$, as a probability distribution over actions such that $\pi\in\triangle(\mathcal{A})$. This implies that $\sum_{a=1}^{A}\pi(a)=1$.
Define $[n]=\{1,2,\ldots,n\}$. In a multi-task structured bandit setting the expected reward for each action in each task is assumed to be an unknown function of the hidden parameter and action features (Lattimore and Szepesvári, 2020; Gupta et al., 2020). The interaction proceeds iteratively over $n$ rounds for each task $m\in[M]$. At each round $t\in[n]$ for each task $m\in[M]$, the learner selects an action $I_{m,t}\in\mathcal{A}$ and observes the reward $r_{m,t}=f(\mathbf{x}(I_{m,t}),\bm{\theta}_{m,*})+\eta_{m,t}$, where $\bm{\theta}_{m,*}\in\mathbb{R}^{d}$ is the hidden parameter specific to the task $m$ to be learned by the learner. The function $f(\cdot,\cdot)$ is the unknown reward structure. This can be $f(\mathbf{x}(I_{m,t}),\bm{\theta}_{m,*})=\mathbf{x}(I_{m,t})^{\top}\bm{\theta}_{m,*}$ for the linear setting or even more complex correlation between features and $\bm{\theta}_{m,*}$ (Filippi et al., 2010; Abbasi-Yadkori et al., 2011; Riquelme et al., 2018; Lattimore and Szepesvári, 2019; Dong et al., 2021). We further assume that the noise $\eta_{m,t}$ is $\sigma^{2}$ sub-Gaussian.
In our paper, we assume that there exist weak demonstrators denoted by $\pi^{w}$. These weak demonstrators are stochastic $A$-armed bandit algorithms like Upper Confidence Bound (UCB) (Auer et al., 2002; Auer and Ortner, 2010) or Thompson Sampling (Thompson, 1933; Agrawal and Goyal, 2012; Russo et al., 2018; Zhu and Tan, 2020). The $\pi^{w}$ at every round $t$ for a task $m\in[M]$ samples an action $I_{m,t}$ and observes the reward $r_{m,t}=f(\mathbf{x}(I_{m,t}),\bm{\theta}_{m,*})+\eta_{m,t}$. Crucially note that $\pi^{w}$ does not observe the feature $\mathbf{x}(I_{m,t})$. Moreover, we assume that the $\pi^{w}$ does not have access to the feature space $\mathcal{X}$ for any task $m$, and hence cannot plan its sampling policy based on the feature vectors. The $\pi^{w}$ is used to collect data on which we train our model. So we call these algorithms weak demonstrators. Note that in many real-world world settings, like robotics (Fong et al., 2003; Lauri et al., 2022), medical treatment (Murphy, 2003) the task is only partially observed or have a complex correlated structure and therefore a learning agent may not have access to the full set of action features, but only observe the reward and action history. This is characterized by $\pi^{w}$. In the next section, we propose the learning model for the training procedure.
@@ 2.2 In-Context Learning Model
Similar to Lee et al. (2023); Sinii et al. (2023); Lin et al. (2023); Ma et al. (2023); Liu et al. (2023c, a) we assume the in-context learning model. We first discuss the pretraining procedure.
Pretraining: Let $\mathcal{T}_{\text{pre}}$ denote the distribution over tasks $m$ at the time of pretraining. Note that by our definition of the task $m$, the distribution $\mathcal{T}_{\text{pre }}$ can span different reward and action spaces. Let $\mathcal{D}_{\text{pre}}$ be the distribution over all possible interactions that the $\pi^{w}$ can generate. We first sample a task $m\sim\mathcal{T}_{\text{pre }}$ and then a context (or a prompt) $\mathcal{H}_{m}$ which is a sequence of interactions for $n$ rounds conditioned on the task $m$ such that $\mathcal{H}_{m}\sim\mathcal{D}_{\text{pre}}(\cdot|m)$. So $\mathcal{H}_{m}=\{I_{m,t},r_{m,t}\}_{t=1}^{n}$. We call this dataset $\mathcal{H}_{m}$ an in-context dataset as it contains the contextual information about the task $m$. We denote the samples in $\mathcal{H}_{m}$ till round $t$ as $\mathcal{H}_{m}^{t}=\{I_{m,s},r_{m,s}\}_{s=1}^{t-1}$. This dataset $\mathcal{H}_{m}$ can be collected in several ways: (1) random interactions within $m$, (2) demonstrations from an expert, and (3) rollouts of an algorithm. Finally, we train a causal GPT-2 transformer model $\mathrm{TF}$ parameterized by $\mathbf{\Theta}$ on this dataset $\mathcal{D}_{\text{pre}}$. Specifically, we define $\mathrm{TF}_{\mathbf{\Theta}}\left(\cdot\mid\mathcal{H}^{t}_{m}\right)$ as the transformer model that observes the dataset $\mathcal{H}^{t}_{m}$ till round $t$ and then produces a distribution over the actions. Our primary novelty lies in our training procedure which we explain in detail in Section 3.1.
Testing: We now discuss the testing procedure for our setting. Let $\mathcal{T}_{\text{test}}$ denote the distribution over test tasks $m\in[M_{\text{test }}]$ at the time of testing. Let $\mathcal{D}_{\text{test}}$ denote a distribution over all possible interactions that can be generated by $\pi^{w}$ during test time. For online training, the dataset $\mathcal{H}_{m}=\{\emptyset\}$ is initialized empty. At each round $t$, $\mathrm{TF}_{\mathbf{\Theta}}(\cdot\mid\mathcal{H}^{t}_{m})$ is deployed where the model samples $I_{t}\sim\mathrm{TF}_{\mathbf{\Theta}}(\cdot\mid\mathcal{H}^{t}_{m})$. The model observes the reward $r_{t}$ which is added to $\mathcal{H}^{t}_{m}$. So the $\mathcal{H}_{m}$ during online testing consists of $\{I_{t},r_{t}\}_{t=1}^{n}$ collected during testing. This interaction procedure is conducted for each test task $m\in[M_{\text{test }}]$. Finally, note that in this testing phase, the model parameter $\mathbf{\Theta}$ is not updated. Finally, the goal of the learner is to minimize cumulative regret for all task $m\in[M_{\text{test}}]$ defined as follows: $\mathbb{E}[R_{n}]=\frac{1}{M_{\text{test}}}\sum_{m=1}^{M_{\text{test}}}\sum_{t=1}^{n}\max_{a\in\mathcal{A}}f\left(\mathbf{x}(m,a)-\mathbf{x}(m,I_{t}),\bm{\theta}_{m,*}\right)$.
@@ 2.3 Related In-context Learning Algorithms
$\mathcal{L}^{\mathrm{AD}}_{t}=\operatorname{cross-entropy}(\mathrm{TF}_{\mathbf{\Theta}}(\cdot|\mathcal{H}_{m}^{t}),p({I}_{m,t}))$ (2)
where $p({I}_{m,t})$ is a one-hot vector such that $p(j)=1$ when $j={I}_{m,t}$ (the true action taken by the demonstrator) and 00 otherwise. During online test time evaluation the AD selects $I_{t}\sim\mathrm{softmax}^{\tau}_{a}(\mathrm{TF}_{\mathbf{\Theta}}(\cdot|\mathcal{H}^{t}_{m}))$. Unfortunately, AD will, at best, match the regret of the demonstrator and cannot obtain lower regret. In the next section, we introduce a new method that can improve upon the demonstrator without knowledge of the optimal action.

### v2

@@ Abstract
In this paper, we study the multi-task structured bandit problem where the goal is to learn a near-optimal algorithm that minimizes cumulative regret. The tasks share a common structure and any optimal algorithm should exploit the shared structure to minimize the cumulative regret for an unseen but related test task. We use a transformer as a decision-making algorithm to learn this shared structure so as to generalize to the unseen test task. The prior work of pretrained decision transformers like DPT requires access to the optimal action during training which may be hard in several scenarios. Diverging from these works, our learning algorithm does not need the knowledge of optimal action per task during training but predicts a reward vector for each of the actions using only the observed offline data from the diverse training tasks. Finally, during inference time, it selects action using the reward predictions employing various exploration strategies in-context for an unseen test task. We show that our model outperforms other methods like DPT, and Algorithmic Distillation (AD) and matches the performance of algorithms that requires privileged information on the structure of the problem. Interestingly, we show that our algorithm, without the knowledge of the underlying problem structure, can learn a near-optimal policy in-context by leveraging the shared structure across diverse tasks. We show that when the shared structure breaks down with the introduction of new actions both during training and test time, our proposed algorithm fails to learn the underlying latent structure. We further show that our algorithm conducts an implicit two-phase exploration and validate all of these findings over several experiments spanning linear, non-linear, real-life datasets, bilinear, and latent bandit settings. Finally, we theoretically analyze the performance of our algorithm and obtain generalization bounds in the in-context multi-task learning setting.
@@ 1 Introduction
In this paper, we study multi-task bandit learning with the goal of learning an algorithm that discovers and exploits structure in a family of related tasks. In multi-task bandit learning, we have multiple distinct bandit tasks for which we want to learn a policy. Though distinct, the tasks share some structure, which we hope to leverage to speed up learning on new instances in this task family. Traditionally, the study of such structured bandit problems has relied on knowledge of the problem structure like linear bandits (Li et al., 2010; Abbasi-Yadkori et al., 2011; Degenne et al., 2020), bilinear bandits (Jun et al., 2019), hierarchical bandits (Hong et al., 2022a; b), Lipschitz bandits (Bubeck et al., 2008; 2011; Magureanu et al., 2014), other structured bandits settings (Riquelme et al., 2018; Lattimore & Szepesvári, 2019; Dong et al., 2021) and even linear and bilinear multi-task bandit settings (Yang et al., 2022a; Du et al., 2023; Mukherjee et al., 2023). When structure is unknown an alternative is to adopt sophisticated model classes, such as kernel machines or neural networks, exemplified by kernel or neural bandits (Valko et al., 2013; Chowdhury & Gopalan, 2017; Zhou et al., 2020; Dai et al., 2022). However, these approaches are also costly as they learn complex, nonlinear models from the ground up without any prior data (Justus et al., 2018; Zambaldi et al., 2018).
@@
In this paper, we consider an alternative approach of synthesizing a bandit algorithm from historical data where the data comes from recorded bandit interactions with past instances of our target task family. Concretely, we are given a set of state-action-reward tuples obtained by running some bandit algorithm in various instances from the task family. We then aim to train a transformer (Vaswani et al., 2017) from this data such that it can learn in-context to solve new task instances. Laskin et al. (2022) consider a similar goal and introduce the Algorithm Distillation (AD) method, however, AD aims to copy the algorithm used in the historical data and thus is limited by the ability of the data collection algorithm. Lee et al. (2023) develop an approach, DPT, that enables learning a transformer that obtains lower regret in-context bandit learning compared to the algorithm used to produce the historical data. However, this approach requires knowledge of the optimal action at each stage of the decision process. In real problems, this assumption is hard to satisfy and we will show that DPT performs poorly when the optimal action is only approximately known. With this past work in mind, the goal of this paper is to answer the question:
@@
Can we learn an in-context bandit learning algorithm that obtains lower regret than the algorithm used to produce the training data without knowledge of the optimal action in each training task?
@@
To answer this question, we introduce a new pre-training methodology, called Pre-trained Decision Transformer with Reward Estimation (PreDeToR) that obviates the need for knowledge of the optimal action in the in-context data –- a piece of information that is often inaccessible. Our key observation is that while the mean rewards of each action change from task to task, certain probabilistic dependencies are persistent across all tasks with a given structure (Yang et al., 2020; 2022a; Mukherjee et al., 2023). These probabilistic dependencies can be learned from the pretraining data and exploited to better estimate mean rewards and improve performance in a new unknown test task. The nature of the probabilistic dependencies depends on the specific structure of the bandit and can be complex (i.e., higher-order dependencies beyond simple correlations). We propose to use transformer models as a general-purpose architecture to capture the unknown dependencies by training transformers to predict the mean rewards in each of the given trajectories (Mirchandani et al., 2023; Zhao et al., 2023). The key idea is that transformers have the capacity to discover and exploit complex dependencies in order to predict the rewards of all possible actions in each task from a small history of action-reward pairs in a new task. This paper demonstrates how such an approach can achieve lower regret by outperforming state-of-the-art baselines, relying solely on historical data, without the need for any supplementary information like the action features or knowledge of the complex reward models. We also show that the shared actions across the tasks are vital for PreDeToR to exploit the latent structure. We show that PreDeToR learns to adapt, in-context, to novel actions and new tasks as long as the number of new actions is small compared to shared actions across the tasks.
@@
    We introduce a new pre-training procedure of learning the underlying reward structure and a decision algorithm. Moreover, PreDeToR by predicting the next reward for all arms circumvents the issue of requiring access to the optimal (or approximately optimal) action during training time.
@@
    We demonstrate empirically that this training procedure results in lower regret in a wide series of tasks (such as linear, nonlinear, bilinear, and latent bandits) compared to prior in-context learning algorithms and bandit algorithms with privileged knowledge of the common structure.
@@
    We also show that our training procedure leverages the shared latent structure. We systematically show that when the shared structure breaks down no reward structure or exploration is learned.
@@ 2 Background
In this section, we first introduce our notation and the multi-task, structured bandit setting. We then formalize the in-context bandit learning model studied in Laskin et al. (2022); Lee et al. (2023); Sinii et al. (2023); Lin et al. (2023); Ma et al. (2023); Liu et al. (2023c; a).
@@ 2.1 Preliminaries
In this paper, we consider the multi-task linear bandit setting (Du et al., 2023; Yang et al., 2020; 2022a). In the multi-task setting, we have a family of related bandit problems that share an action set $\mathcal{A}$ and also a common action feature space $\mathcal{X}$. The actions in $\mathcal{A}$ are indexed by $a=1,2,\ldots,A$. The feature of each action is denoted by $\mathbf{x}(a)\in\mathbb{R}^{d}$ and $d\ll A$. A policy, $\pi$, is a probability distribution over the actions.
Define $[n]=\{1,2,\ldots,n\}$. In a multi-task structured bandit setting the expected reward for each action in each task is assumed to be an unknown function of the hidden parameter and action features (Lattimore & Szepesvári, 2020; Gupta et al., 2020). The interaction proceeds iteratively over $n$ rounds for each task $m\in[M]$. At each round $t\in[n]$ for each task $m\in[M]$, the learner selects an action $I_{m,t}\in\mathcal{A}$ and observes the reward $r_{m,t}=f(\mathbf{x}(I_{m,t}),\bm{\theta}_{m,*})+\eta_{m,t}$, where $\bm{\theta}_{m,*}\in\mathbb{R}^{d}$ is the hidden parameter specific to the task $m$ to be learned by the learner. The function $f(\cdot,\cdot)$ is the unknown reward structure. This can be $f(\mathbf{x}(I_{m,t}),\bm{\theta}_{m,*})=\mathbf{x}(I_{m,t})^{\top}\bm{\theta}_{m,*}$ for the linear setting or even more complex correlation between features and $\bm{\theta}_{m,*}$ (Filippi et al., 2010; Abbasi-Yadkori et al., 2011; Riquelme et al., 2018; Lattimore & Szepesvári, 2019; Dong et al., 2021).
In our paper, we assume that there exist weak demonstrators denoted by $\pi^{w}$. These weak demonstrators are stochastic $A$-armed bandit algorithms like Upper Confidence Bound (UCB) (Auer et al., 2002; Auer & Ortner, 2010) or Thompson Sampling (Thompson, 1933; Agrawal & Goyal, 2012; Russo et al., 2018; Zhu & Tan, 2020). We refer to these algorithms as weak demonstrators because they do not use knowledge of task structure or arm feature vectors to plan their sampling policy. In contrast to a weak demonstrator, a strong demonstrator, like LinUCB, uses feature vectors and knowledge of task structure to conduct informative exploration. Whereas weak demonstrators always exist, there are many real-world settings with no known strong demonstrator algorithm or where the feature vectors are unobserved and the learner can only use the history of rewards and actions.
@@ 2.2 In-Context Learning Model
Similar to Lee et al. (2023); Sinii et al. (2023); Lin et al. (2023); Ma et al. (2023); Liu et al. (2023c; a) we assume the in-context learning model. We first discuss the pretraining procedure.
Pretraining: Let $\mathcal{T}_{\text{pre}}$ denote the distribution over tasks $m$ at the time of pretraining. Let $\mathcal{D}_{\text{pre}}$ be the distribution over all possible interactions that the $\pi^{w}$ can generate. We first sample a task $m\sim\mathcal{T}_{\text{pre }}$ and then a context $\mathcal{H}_{m}$ which is a sequence of interactions for $n$ rounds conditioned on the task $m$ such that $\mathcal{H}_{m}\sim\mathcal{D}_{\text{pre}}(\cdot|m)$. So $\mathcal{H}_{m}=\{I_{m,t},r_{m,t}\}_{t=1}^{n}$. We call this dataset $\mathcal{H}_{m}$ an in-context dataset as it contains the contextual information about the task $m$. We denote the samples in $\mathcal{H}_{m}$ till round $t$ as $\mathcal{H}_{m}^{t}=\{I_{m,s},r_{m,s}\}_{s=1}^{t-1}$. This dataset $\mathcal{H}_{m}$ can be collected in several ways: (1) random interactions within $m$, (2) demonstrations from an expert, and (3) rollouts of an algorithm. Finally, we train a causal GPT-2 transformer model $\mathrm{TF}$ parameterized by $\mathbf{\Theta}$ on this dataset $\mathcal{D}_{\text{pre}}$. Specifically, we define $\mathrm{TF}_{\mathbf{\Theta}}\left(\cdot\mid\mathcal{H}^{t}_{m}\right)$ as the transformer model that observes the dataset $\mathcal{H}^{t}_{m}$ till round $t$ and then produces a distribution over the actions. Our primary novelty lies in our training procedure which we explain in detail in Section 3.1.
Testing: We now discuss the testing procedure for our setting. Let $\mathcal{T}_{\text{test}}$ denote the distribution over test tasks $m\in[M_{\text{test }}]$ at the time of testing. Let $\mathcal{D}_{\text{test}}$ denote a distribution over all possible interactions that can be generated by $\pi^{w}$ during test time. At deployment time, the dataset $\mathcal{H}_{m}^{0}\leftarrow\{\emptyset\}$ is initialized empty. At each round $t$, an action is sampled from the trained transformer model $I_{t}\sim\mathrm{TF}_{\mathbf{\Theta}}(\cdot\mid\mathcal{H}^{t}_{m})$. The sampled action and resulting reward, $r_{t}$, are then added to $\mathcal{H}^{t}_{m}$ to form $\mathcal{H}^{t+1}_{m}$ and the process repeats for $n$ total rounds. Finally, note that in this testing phase, the model parameter $\mathbf{\Theta}$ is not updated. Finally, the goal of the learner is to minimize cumulative regret for all task $m\in[M_{\text{test}}]$ defined as follows: $\mathbb{E}[R_{n}]=\frac{1}{M_{\text{test}}}\sum_{m=1}^{M_{\text{test}}}\sum_{t=1}^{n}\max_{a\in\mathcal{A}}f\left(\mathbf{x}(a),\bm{\theta}_{m,*}\right)-f\left(\mathbf{x}(I_{t}),\bm{\theta}_{m,*}\right)$.
@@ 2.3 Related In-context Learning Algorithms
$\mathcal{L}^{\mathrm{{\color[rgb]{0,0.6,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.6,0}AD}}}_{t}=\operatorname{cross-entropy}(\mathrm{TF}_{\mathbf{\Theta}}(\cdot|\mathcal{H}_{m}^{t}),p({I}_{m,t}))$ (2)
where $p({I}_{m,t})$ is a one-hot vector such that $p(j)=1$ when $j={I}_{m,t}$ (the true action taken by the demonstrator) and 00 otherwise. At deployment time, AD selects $I_{t}\sim\mathrm{softmax}^{\tau}_{a}(\mathrm{TF}_{\mathbf{\Theta}}(\cdot|\mathcal{H}^{t}_{m}))$. The objective of AD is to match the performance of the demonstrator. In the next section, we introduce a new method that can improve upon the demonstrator without knowledge of the optimal action.